For a given gene locus, you identify the following genotypic frequencies: AA = 0.65 BB = 0.25

What is the frequency of the AB genotype?

What is the frequency of the A allele?

What is the frequency of the B allele?

Show all work.
Answer #1

Given genotypic frequencies are

AA= 0.65

BB = 0.25

According to HW equations, the genotypic frequencies of all alleles is equal to 1.

P^2 = 0.65

q^2 = 0.25

2pq is frequency of AB.

That means 0.65+0.25+ frequency of AB = 1

Therefore frequency of AB genotype = 1-(0.65+0.25) = 1-0.90 = 0.10.

Frequency of A allele is calculate by taking the number of alleles of each type and dividing it by total number of alleles.

Now, AA frequency is 0.65, that means 65 in 100. It has two A alleles, so total A alleles = 65*2=130.

The heterozygote AB also has one allele. There are 10 AB (0.1) so total of A alleles = 130+10=140.

Total individuals are 100, so total alleles will be 200.

Hence frequency of A allele = 140/200= 0.7

Similarly, the frequency of B allele = BB = 25*2=50; AB = 10.

Frequency of B allele = 60/200=0.3.